# Stronger Language Models Produce More Human-Like Errors

### v1

@@ Abstract
Do language models converge toward human-like reasoning patterns as they improve? We provide surprising evidence that while overall reasoning capabilities increase with model sophistication, the nature of errors increasingly mirrors predictable human reasoning fallacies: a previously unobserved inverse scaling phenomenon. To investigate this question, we apply the Erotetic Theory of Reasoning (ETR), a formal cognitive framework with empirical support for predicting human reasoning outcomes. Using the open-source package PyETR, we generate logical reasoning problems where humans predictably err, evaluating responses from 38 language models across 383 reasoning tasks. Our analysis indicates that as models advance in general capability (as measured by Chatbot Arena scores), the proportion of their incorrect answers that align with ETR-predicted human fallacies tends to increase ($\rho=0.360,p=0.0265$). Notably, as we observe no correlation between model sophistication and logical correctness on these tasks, this shift in error patterns toward human-likeness occurs independently of error rate. These findings challenge the prevailing view that scaling language models naturally obtains normative rationality, suggesting instead a convergence toward human-like cognition inclusive of our characteristic biases and limitations, as we further confirm by demonstrating order-effects in language model reasoning.
@@ 1 Introduction
As language models rapidly advance in capabilities, they increasingly demonstrate sophisticated reasoning abilities across diverse domains [1, 2]. A naive prediction would have future language models achieving perfect rationality simply through scale. However, as these models become more capable of mimicking human language and problem-solving, do they also inherit our systematic reasoning errors?
Human reasoning exhibits robust patterns of fallacious inference that have been extensively documented in cognitive science [3, 4, 5, 6, 7]. These patterns are not random mistakes, but systematic deviations from normative standards that persist even among educated participants and that follow predictable patterns across diverse reasoning tasks. The study of these fallacies has been central to our understanding of human cognition and has profound implications for fields ranging from economics to medical decision-making.
Our work investigates whether increasingly sophisticated language models show a similar pattern of fallacious reasoning to humans. This question has both theoretical and practical significance. Theoretically, it provides insight into whether human-like reasoning errors emerge naturally from systems trained on human-generated text. Practically, it raises important considerations for AI systems deployed in contexts requiring reliable reasoning, such as medical diagnosis, legal analysis, or high-stakes decision support [8, 9].
@@
ETR has been implemented as PyETR [13], an open-source Python library. Using PyETR, we mechanically generate reasoning problems predicted to lead human-like reasoners into making errors. For our experiments we fix a diverse corpus of 383 such problems for which ETR predicts a specific fallacious response.
Our experimental investigation evaluates 38 language models ranging from smaller models (e.g., Mistral 7B Instruct v0.1) to state-of-the-art systems (e.g., GPT-4.5, Claude 3.7), using Chatbot Arena Elo as a proxy for model sophistication [14]. Our central finding is striking: as model capability increases, models become significantly more likely to commit human-like reasoning errors ($\rho=0.360,p=0.0265$).
This finding represents a notable example of an inverse scaling law in AI development—where model improvement on general metrics corresponds to decreased performance on specific desirable properties. While language models have demonstrated improved performance with scale on many reasoning tasks [1, 15], our work suggests that they may simultaneously become more susceptible to human-like reasoning fallacies. Beyond practical implications, our findings provide novel evidence relevant to the debate on whether LLMs develop internal reasoning processes analogous to human cognition, or simply mimic surface-level patterns from their training data.
@@
(1) We empirically demonstrate a significant inverse scaling relationship between language model sophistication and susceptibility to human-like reasoning fallacies. (2) We provide a theoretical analysis connecting our findings to broader questions about the emergence of human-like cognitive properties in language models. (3) With our methodology, we give the first substantial application of PyETR, an open-source computational implementation of the Erotetic Theory of Reasoning, to generate reasoning problems with predictable human-like failure patterns.
The remainder of this paper is organized as follows: Section 2 introduces the Erotetic Theory of Reasoning and our PyETR implementation. Section 3 details our experimental methodology. Section 4 presents our results. Section 5 discusses and concludes.
@@ 2 The Erotetic Theory of Reason
Modern understanding of human reasoning reveals a fascinating duality: remarkable achievements coexist with persistent, predictable errors. The Erotetic Theory of Reasoning (ETR) provides a unified framework explaining both capacities through a single mechanism: humans construe problems as questions requiring answers, mentally representing disjunctive alternatives that they filter based on how well they match new information. This filtering process, where we prematurely dispose of alternatives before incorporating new data, enables cognitive efficiency but introduces systematic biases. Importantly, ETR demonstrates that reasoning success or failure depends not on separate cognitive systems but on whether one achieves “erotetic equilibrium,” a state where conclusions remain stable against potential follow-up questions.
At its core, ETR (full formalism in Appendix A) proposes a computational model where reasoning operates through three key mechanisms: (1) maintaining disjunctive alternatives representing possible answers to implicit questions, (2) filtering these alternatives by best-match with new information, potentially eliminating relevant possibilities prematurely, and (3) correcting potential errors by raising explicit questions that reintroduce necessary alternatives. This conceptually parsimonious framework has demonstrated remarkable empirical success in predicting systematic reasoning fallacies across propositional and first-order logic, probability theory, and decision-making. Mathematical proofs further establish that as reasoners approach complete erotetic equilibrium by raising appropriate questions, their conclusions converge to normative standards including logical soundness and completeness, Bayesian updating, and expected utility maximization, making ETR a compelling explanation for human reasoning as a unified cognitive capacity whose errors arise from computational constraints rather than irrationality.
Let us demonstrate the basics of ETR with the following reasoning problem, Example 49 in [10]. Premise 1 Either there is an ace in Mary’s hand and some other player has a king, or else there is a queen in John’s hand and some other player has a jack. Premise 2 Sally has a king. Question Does it follow that Mary has an ace? While the correct logical answer to the question is “not necessarily”, the conclusion “Mary has an ace” was endorsed by 60% of respondents (unpublished data following the methods of [11]).
@@
To model the problem with logic, we would first choose a representation of the basic predicates, in this case $\mathsf{Ace}(x)$ for the predicate that (the card) $x$ is an ace, and $\mathsf{Has}(y,z)$ for the predicate that (player) $y$ holds (a card) $z$. The same applies when using ETR, where instead of logical formulas we have the central concept of view (see Appendix A) for the full mathematical formalism). It will suffice for now to use the (still precise) shorthand notation to express the premises and conclusion as ETR views: [^1] An additional aspect not present in ordinary logic is issue structure. This is notated with an asterisk on some occurrences of terms, indicative of being ‘at issue’ for the containing atomic proposition. The issue structure has no logical content, but plays a role in guiding ETR inferences. Determination of issue structure from cues in natural language or context is outside the scope of ETR, but a simple heuristic in this example is that the repetition of ‘Ace’ and ‘King’ suggests those concepts are at issue. P1 $\{\mathsf{Ace}(a^{*})\mathsf{Has}(\mathsf{Mary},a)\mathsf{King}(b^{*})\mathsf{Has}(c,b),\mathsf{Queen}(d)\mathsf{Has}(\mathsf{John},d)\mathsf{Jack}(e)\mathsf{Has}(f,e)\}$ P2 $\{\mathsf{King}(g^{*})\mathsf{Has}(\mathsf{Sally},g)\}$ Q $\{\mathsf{Ace}(h^{*})\mathsf{Has}(\mathsf{Mary},h)\}$
@@
Each view is given as a disjunctive set of alternatives, separated by a comma. Each alternative is to be considered as a conjunctive set of atomic propositions, where for convenience we omit the set-brackets and commas and simply juxtapose elements. Many readers will recognize these as disjunctive normal forms, but in ETR there is no inherent ordering to the atoms in a conjunction or alternatives in the disjunction. The individual lowercase letters can be understood are as existentially quantified variables, with the scope of the quantification encompassing the entire disjunction (‘prenex form’). A view may also carry a supposition, written as a superscript (see Table 1 for examples).
@@
$>$ from pyetr.inference import default_procedure_does_it_follow
$>$ default_procedure_does_it_follow([p1,p2],cc)
@@ 3 Methodology
In our experiment, a reasoning problem is given by a list of views (the premises), and the problem is to answer the question “what if anything follows?”. We are most interested in reasoning problems where the ETR-predicted answer to this question is a logical fallacy.
@@ 3.3 Mapping to natural language
To evaluate language model reasoning in a more naturalistic context, we developed a systematic approach to convert formal logical statements into natural language prompts. We created 12 themes for natural language framings of logical problems, such as a researcher figuring out the properties of a novel element or a newly discovered creature. For each theme, we established consistent mappings between logical elements and thematic equivalents. The thematic conversion process serves multiple purposes: it ensures our results are robust against content effects [24]; it enables us to test whether the reasoning patterns are consistent in diverse contexts; and it mitigates potential data contamination, a significant problem for LLM evaluation [30, 31, 32, 33] by reformulating problems in novel scenarios unlikely to appear in training data.
@@
This exact prompt was sent to the LLM. It includes a thematic preamble and a natural language prompt to reason in an open-ended fashion with: “What if anything follows?” Each problem maintained the same prompt structure: an introduction establishing the thematic context, followed by the premises expressed in natural language, and concluding with a standardized question asking what follows from the premises. The prompt explicitly stated that nothing might follow, to avoid biasing models toward making unwarranted inferences [25].
@@ 3.4 Data Collection
We evaluated 38 language models of varying sophistication, including several specifically designed as reasoning models. Selection informally prioritised quantity, popularity of usage, and representation of a diverse range of scores. All models were accessed via OpenRouter and tested using LLMharness. Chatbot Arena scores were obtained from https://huggingface.co/spaces/lmarena-ai/chatbot-arena-leaderboard/resolve/main/elo_results_20250423.pkl, released under the Apache 2.0 license. To ensure a fair comparison across architectural differences, all reasoning models were allocated 2,400 thinking tokens during evaluation. This standardization was justified given our use of the Chatbot Arena Elo score as our primary capability metric, which already accounts for performance differences regardless of underlying mechanisms [14].
@@ 4 Experimental Results
We operationalize a “human-like fallacy” (or simply “fallacy”) as a response that is both ETR-predicted and logically incorrect. This definition captures instances where models make the same systematic reasoning errors that ETR predicts humans would make. Formally, for a given model $m$ and reasoning problem $p$, we define:
[…]
@@ 5 Conclusion and Discussion
Interpretation of Key Findings Our results reveal a statistically significant trend: as language models advance in capability (measured by Chatbot Arena Elo), their logical errors increasingly align with human errors predicted by ETR. Critically, while stronger language models demonstrate improved capabilities across most benchmarks, we observe a systematic increase in human-like fallibility: an inverse scaling trend that persists across model families, architectures, and training paradigms.
Moreover, we demonstrate order-effects for LLM reasoning analogous to those observed in humans. This finding suggests that beyond merely emulating modal human reasoning patterns, LLMs exhibit signs of mirroring human cognition at deeper levels, including characteristic fallacies.
@@
Theoretical Implications These findings challenge the assumption that scaling alone leads to more rational systems. Instead, we observe convergence toward human-like reasoning patterns including characteristic fallacies, suggesting that current language models may be developing internal processes that functionally mirror human cognitive shortcuts rather than idealized rational agents. If models increasingly adopt human-like reasoning patterns, including our characteristic errors, this suggests the emergence of cognitive structures functionally similar to human reasoning, rather than alternative solution methods that happen to produce similar outputs.
Connection to Inverse Scaling Laws and Practical Applications Our findings represent an interesting case of an inverse scaling law [45], where improvement on general metrics correlates with increasing susceptibility to specific reasoning fallacies. This phenomenon has both theoretical significance and practical applications for AI development.
Our results provide a foundation for several promising research directions. For benchmark development, our approach offers a systematic method for generating reasoning problems where human-like fallacies are predicted, creating evaluation sets that could complement existing reasoning benchmarks. For capabilities, the identified fallacy patterns could serve as targeted training examples to fortify models against human-like reasoning errors while maintaining their general capabilities. As a tool for assessing engineering choices, our methodology provides a quantitative framework for evaluating the effectiveness of various interventions aimed at improving model reasoning, from prompt engineering to architectural modifications.
Broader Implications This inverse scaling phenomenon has profound implications for AI alignment. As models become more capable, anticipating and mitigating human-like reasoning errors becomes increasingly important, especially in high-stakes contexts requiring reliable reasoning, such as medical diagnosis, legal analysis, or decision support [8, 9].
We see these results as opening new avenues for understanding both artificial and human reasoning. As language models continue to advance, this work provides a foundation for developing systems that combine human-like understanding with more robust reasoning capabilities: systems that can successfully reason as we do while avoiding the pitfalls that characterize human cognition.
@@ Appendix A A formal presentation of the Erotetic Theory
[…]
@@
?$\in\mathbf{T}_{1}$ and $\langle f^{k},\langle t_{1},\ldots,t_{k}\rangle\rangle\in\mathbf{T}_{1}$ whenever $f^{k}\in\mathbf{F}$ and for some $1\leq i\leq k$, $t_{i}\in\mathbf{T}_{1}$ and $t_{j}\in\mathbf{T}$ for $j\neq i$. $\langle P^{k},\langle t_{1},\ldots,t_{k}\rangle\rangle\in\mathcal{A}_{1}$, whenever $P^{k}\in\mathbf{P}$ and for some $1\leq i\leq k$, $t_{j}\in\mathbf{T}$ for all $j\neq i$ and $t_{i}\in\mathbf{T}_{1}$.
@@
$[R]_{X}=\langle U_{R}\cap X,E_{R}\cap X,D_{R}\cap((E_{R}\cap X)\times(U_{R}\cap X))\rangle$.
$[R]_{\Gamma}=[R]_{\mathbf{A}(\Gamma)}$.
@@
$[I]=\{\langle t,x\rangle:\langle t,x\rangle\in I\wedge x[t/?]\in\mathcal{A}(\Gamma\cup\Theta)\}$
@@
$\Gamma^{\Theta}_{RI}[\Delta^{\{0\}}_{SJ}]^{A}=(\operatorname{argmax}\limits_{\gamma\in\Gamma}\ \Delta[\{\{p\}:p\in\gamma\}]^{\mathcal{A}P})^{\Theta}_{[R][I]}$
@@
$[\Gamma^{\Theta}_{RI}]^{N}=(\Theta^{\{0\}}\otimes([\Gamma]^{N})^{\{0\}})_{[R]^{N}[I]^{N}}$
$[\Gamma]^{N}=\bigotimes_{\gamma\in\Gamma}\hat{\gamma}^{\{0\}}$
@@
$\bar{p}=\bar{F}\vec{\tau}\text{ if }p=F\vec{\tau};$ $\bar{p}=F\vec{\tau}\text{ if }p=\bar{F}\vec{\tau}$
$[R]^{N}=\langle E_{R},U_{R},\{\langle a,b\rangle\in U_{R}\times E_{R}:\langle b,a\rangle\notin D_{R}\}\rangle$
$[I]^{N}=I\cup\{\langle t,\bar{x}\rangle:\langle t,x\rangle\in I\}$
@@
$a\in\mathbf{A}$ is novel during an inference step $\langle C,G\rangle[D]^{O}$ if $a\notin\mathbf{A}(G)\cup\mathbf{A}(D)$ and $a$ has not yet appeared in the computation of the conclusion of the step. A function $\nu:X\to\mathbf{A}$, where $X\subseteq\mathbf{A}$ is finite, is novel during an inference step if each $\nu(x)$ is novel and $\nu$ is injective, and then $[\nu]_{X}$ denotes the simultaneous substitution $[\nu(a_{1})/a_{1},\ldots,\nu(a_{k})/a_{k}]$, where $X=\{a_{1},\ldots,a_{k}\}$. Let $(-)^{\mathrm{nov}(X)}$ stand for $(-)[\nu]_{X}$ where $\nu:X\to\mathbf{A}$ is novel.
@@
For either $\mathbf{A}(R)\cap\mathbf{A}(S)=\emptyset$ or $[R]_{\Delta\cup\Psi}=S$,
$\Gamma^{\Theta}_{RI}[\Delta^{\Psi}_{SJ}]^{M}=\bigoplus^{R\bowtie S}_{\gamma\in\Gamma}\Big{(}\{\gamma\}^{\Theta}_{RI}\otimes^{R\bowtie S}\Delta^{\Psi}_{SJ}\otimes^{R\bowtie S}\bigotimes^{R\bowtie S}_{\langle t,u\rangle\in M^{\prime}_{IJ}(\gamma)}\mathrm{Sub}^{R\bowtie S}_{\langle t,u\rangle}(\Delta^{\{0\}}_{J})\Big{)}$,
@@
For $\mathbf{A}(\Gamma)\cap\mathbf{A}(\Theta)=\emptyset$ and either $\mathbf{A}(R)\cap\mathbf{A}(S)=\emptyset$ or $[R]_{\Delta}=S$,
$\Gamma^{\Theta}_{RI}[\Delta^{\{0\}}_{SJ}]^{U}=\{0\}^{\Theta}_{RI}\otimes^{R\bowtie S}\bigotimes^{R\bowtie S}_{\langle u,t\rangle\in M^{\prime}_{IJ}}\mathrm{Sub}^{R\bowtie S}_{\langle t,u\rangle}(\Gamma^{\{0\}}_{I})$, where
@@
    $\mathbf{A}(\Gamma)\cap\mathbf{A}(\Theta)=\emptyset$ and either $\mathbf{A}(R)\cap\mathbf{A}(S)=\emptyset$ or $[R]_{\Delta}=S$.
@@
($(\{0:\neg\exists\delta_{\in\Delta}\exists\gamma.\Phi(\gamma,\delta)\}\cup\{\delta\in\Delta:\exists\gamma_{\in\Gamma}.\Phi(\gamma,\delta)\})^{\Theta}_{[R\bowtie\langle U_{R},E_{S}\backslash E_{R},D_{S^{\prime}}\rangle][I\cup J]}$
@@
(I) If $\mathbf{A}(\Delta)\subseteq\mathbf{A}(\Gamma\cup\Theta)$ and $[R]_{\Delta}=S$,

### v2

# Theory-Grounded Evaluation of Human-Like Fallacy Patterns in LLM Reasoning

@@ Abstract
We study logical reasoning in language models by asking whether their errors follow established human fallacy patterns. Using the Erotetic Theory of Reasoning (ETR) and its open‑source implementation, PyETR, we programmatically generate 383 formally specified reasoning problems and evaluate 38 models. For each response, we judge logical correctness and, when incorrect, whether it matches an ETR‑predicted fallacy. Two results stand out: (i) as a capability proxy (Chatbot Arena Elo) increases, a larger share of a model’s incorrect answers are ETR‑predicted fallacies ($\rho=0.360,p=0.0265$), while overall correctness on this dataset shows no correlation with capability; (ii) reversing premise order significantly reduces fallacy production for many models, mirroring human order effects. Methodologically, PyETR provides an open‑source pipeline for unbounded, synthetic, contamination‑resistant reasoning tests linked to a cognitive theory, enabling analyses that focus on error composition rather than error rate.
@@ 1 Introduction
Language models increasingly solve complex tasks [1, 2]. We ask a simple question: when they err on controlled reasoning problems, do their errors align with human fallacies? Human reasoning shows systematic, repeatable fallacies [3, 4, 5, 6, 7]. These are not random mistakes; they follow predictable patterns across tasks. Understanding whether LLM errors share these patterns is useful both scientifically and for deployment in settings that require reliable reasoning [8, 9].
@@
ETR has an open-source implementation, PyETR [13]. Using PyETR, we mechanically generate reasoning problems predicted to elicit specific fallacies. We evaluate 38 models on a fixed corpus of 383 such problems.
Our experimental investigation evaluates 38 language models ranging from smaller models (e.g., Mistral 7B Instruct v0.1) to larger systems (e.g., GPT-4.5, Claude 3.7), treating Chatbot Arena Elo as one capability proxy [14]. We observe that as this proxy increases, a larger share of logically incorrect answers align with ETR-predicted human fallacies ($\rho=0.360,p=0.0265$). We do not claim causality and restrict interpretation to this evaluation setting.
While language models often improve on many benchmarks [1, 15], our results indicate that within our task domain, the proportion of errors matching ETR-predicted fallacies increases with Chatbot Arena Elo as a capability proxy. These findings are consistent with overlap between LLM error patterns and human fallacy patterns and do not require commitments about underlying cognitive mechanisms.
We check the robustness of our result by using metrics other than Chatbot Arena Elo where available for a subset of the 38 models. We found significant Spearman correlation with estimates of training compute [16] ($\rho=0.489,p=0.0334$) and significant exponential fit for mean score on the HELM capabilities benchmark [17] ($r=0.796,p=0.0103$).
@@
We report a statistically significant correlation between Chatbot Arena Elo as a capability proxy and reasoning errors predicted by ETR, an empirically grounded formal theory of human reasoning. We also provide the first substantial application of PyETR — an open-source implementation of ETR — to generate unbounded, synthetic, contamination-resistant reasoning tasks linking model errors to theory-predicted fallacies.
The remainder of the paper is organised as follows. Section 2 introduces ETR and PyETR. Section 3 details methodology. Section 4 presents results. Section 5 discusses implications.
@@ 2 The Erotetic Theory of Reason
Modern accounts of human reasoning highlight a duality: impressive competence and predictable errors. The Erotetic Theory of Reasoning (ETR) explains both with one idea: we reason by managing questions and candidate answers (“alternatives”) and filtering them when new information arrives. Filtering is efficient but can drop relevant alternatives too early, producing characteristic fallacies. ETR formalises this process and defines when conclusions stabilise against follow-up questions (“erotetic equilibrium”).
At a high level (full formalism in Appendix A), ETR has three ingredients: (1) maintain disjunctive alternatives as candidate answers to an implicit question, (2) filter by best match with incoming information (risking premature elimination), and (3) recover by asking questions (raising structured alternatives) that reintroduce necessary information. This predicts a range of empirical reasoning data across propositional and first-order logic, probability, and decision-making. Morover, it is mathematically proved in [10] that ETR converges to normative standards of rationality in the limit, as the recovery step introduces enough alternatives to stabilise further reasoning steps.
Let us demonstrate ETR with a compact example (Example 49 in [10]).
Premise 1 Either there is an ace in Mary’s hand and some other player has a king, or else there is a queen in John’s hand and some other player has a jack.
Premise 2 Sally has a king.
Question Does it follow that Mary has an ace?
The correct logical answer is “not necessarily”, yet many respondents endorse “Mary has an ace” (60% in data following [11]).
@@
To model the problem with logic, we would first choose a representation of the basic predicates, in this case $\mathsf{Ace}(x)$ for the predicate that (the card) $x$ is an ace, and $\mathsf{Has}(y,z)$ for the predicate that (player) $y$ holds (a card) $z$. The same applies when using ETR, where instead of logical formulas we have the central concept of view (see Appendix A) for the full mathematical formalism). It will suffice for now to use the (still precise) shorthand notation to express the premises and conclusion as ETR views: [^1] An additional aspect not present in ordinary logic is issue structure. This is notated with an asterisk on some occurrences of terms, indicative of being ‘at issue’ for the containing atomic proposition. The issue structure has no logical content, but plays a role in guiding ETR inferences. Determination of issue structure from cues in natural language or context is outside the scope of ETR, but a simple heuristic in this example is that the repetition of ‘Ace’ and ‘King’ suggests those concepts are at issue.
P1 $\{\mathsf{Ace}(a^{*})\mathsf{Has}(\mathsf{Mary},a)\mathsf{King}(b^{*})\mathsf{Has}(c,b),\mathsf{Queen}(d)\mathsf{Has}(\mathsf{John},d)\mathsf{Jack}(e)\mathsf{Has}(f,e)\}$
P2 $\{\mathsf{King}(g^{*})\mathsf{Has}(\mathsf{Sally},g)\}$
Q $\{\mathsf{Ace}(h^{*})\mathsf{Has}(\mathsf{Mary},h)\}$
@@
Each view is given as a disjunctive set of alternatives, separated by a comma. Each alternative is to be considered as a conjunctive set of atomic propositions, where for convenience we omit the set-brackets and commas and simply juxtapose elements. Many readers will recognise these as disjunctive normal forms, but in ETR there is no inherent ordering to the atoms in a conjunction or alternatives in the disjunction. The individual lowercase letters can be understood are as existentially quantified variables, with the scope of the quantification encompassing the entire disjunction (‘prenex form’). A view may also carry a supposition, written as a superscript (see Table 1 for examples). Note that since any formula of first-order logic can be written in prenex disjunctive normal form, ETR is as expressive as full first-order logic.
@@
>>> from pyetr.inference import default_procedure_does_it_follow
>>> default_procedure_does_it_follow([p1,p2],$c$)
@@ 3 Methodology
In our experiment, a reasoning problem is given by a list of views (the premises), and the problem is to answer the question “what if anything follows?”. Our method generates reasoning problems where the ETR-predicted answer is a logical fallacy.
@@ 3.1 Data processing
To avoid complications in the mapping to natural language, we restrict to views with only monadic (one-place) predicates. The corresponding monadic first-order logic is known to be less expressive than full first-order logic, but still allows for a rich set of reasoning problems. Our method could easily be extended to include polyadic predicates.
@@ 3.2 Generative pipeline
Dataset curation
We initially generated 400 problems. Pre-analysis integrity checks identified 17 items whose fallacy status did not meet the final specification, so we excluded them prior to all analyses, yielding 383 problems. Given the modest reduction in sample size and the cost of re-running all model evaluations, we report results on this vetted 383-set; the generation pipeline supports regenerating larger sets for future work.
@@ 3.3 Mapping to natural language
To evaluate logical reasoning in language models in a more naturalistic context, we developed a systematic approach to convert formal logical statements into natural language prompts. We created 12 themes for natural language framings of logical problems, such as a researcher figuring out the properties of a novel element or a newly discovered creature. For each theme, we established consistent mappings between logical elements and thematic equivalents. The thematic conversion process serves multiple purposes: it ensures our results are robust against content effects [26]; it enables us to test whether the reasoning patterns are consistent in diverse contexts; and it mitigates potential data contamination, a significant problem for LLM evaluation [32, 33, 34, 35] by reformulating problems in novel scenarios unlikely to appear in training data.
@@
This is typical of the prompts sent to the LLM. It includes a thematic preamble and a natural language prompt to reason in an open-ended fashion with: “What if anything follows?” Each problem maintained the same prompt structure: an introduction establishing the thematic context, followed by the premises expressed in natural language, and concluding with a standardised question asking what follows from the premises. The prompt explicitly stated that nothing might follow, to avoid biasing models toward making unwarranted inferences [27].
@@ 3.4 Data Collection
We evaluated 38 language models of varying sophistication, including several specifically designed as reasoning models. Selection informally prioritized quantity, popularity of usage, and representation of a diverse range of scores. All models were accessed via OpenRouter and tested using LLMharness. Chatbot Arena scores were obtained from https://huggingface.co/spaces/lmarena-ai/chatbot-arena-leaderboard/resolve/main/elo_results_20250423.pkl. Scores for the HELM Capabilities benchmark for 9 models were obtained from https://crfm.stanford.edu/helm/capabilities/v1.10.0/#/leaderboard [17] and estimates for training compute for 19 models were obtained from https://epoch.ai/data/ai-models [16] (both URLs accessed 30 July 2025).
To ensure a fair comparison across architectural differences, all reasoning models were allocated 2,400 thinking tokens during evaluation. This standardization was justified given our use of the Chatbot Arena Elo score as our primary capability metric, which already accounts for performance differences regardless of underlying mechanisms [14].
@@ 4 Experimental Results
We operationalise a “human-like fallacy” (or simply “fallacy”) as a response that is both ETR-predicted and logically incorrect. This definition captures instances where models make the same systematic reasoning errors that ETR predicts humans would make. Formally, for a given model $m$ and reasoning problem $p$, we define:
$\text{HumanLikeFallacy}(m,p)=\begin{cases}1&\text{if ETR-predicted}(m,p)\land\neg\text{LogicallyCorrect}(m,p)\\ 0&\text{otherwise}\end{cases}$
@@ 5 Conclusion and Discussion
Interpretation of Key Findings Our results reveal a statistically significant correlational trend in error composition: as language models advance in capability (proxied by Chatbot Arena Elo), a larger share of logically incorrect answers align with human errors predicted by ETR. While stronger language models demonstrate improved capabilities across many benchmarks, we do not observe increased overall correctness in this dataset. Moreover, we demonstrate order-effects for LLM reasoning analogous to those observed in humans. This indicates overlap between LLM error patterns and established human fallacy patterns on these tasks.
@@
Theoretical Implications These findings challenge the assumption that scaling alone leads to more normatively correct reasoning systems. Instead, within our evaluation setting we observe greater overlap between model error patterns and characteristic human fallacies. This does not by itself imply shared internal processes; it highlights an evaluation axis (error composition) complementary to overall accuracy.
Our results provide a foundation for several research directions. For benchmark development, our approach offers a systematic method for generating reasoning problems where human-like fallacies are predicted, creating evaluation sets that could complement existing reasoning benchmarks. For capabilities, the identified fallacy patterns could serve as targeted training examples to fortify models against human-like reasoning errors while maintaining their general capabilities. As a tool for assessing engineering choices, our methodology provides a quantitative framework for evaluating the effectiveness of various interventions aimed at improving model reasoning, from prompt engineering to architectural modifications.
Broader Implications This observed trade-off in error composition has implications for AI alignment. As models become more capable, anticipating and mitigating human-like reasoning errors becomes increasingly important, especially in high-stakes contexts requiring reliable reasoning, such as medical diagnosis, legal analysis, or decision support [8, 9].
We see these results as opening new avenues for understanding both artificial and human reasoning. As language models continue to advance, this work provides a foundation for developing systems that combine human-like understanding with more robust reasoning capabilities: systems that can successfully reason as we do while avoiding the pitfalls that characterise human cognition.
Author use of AI tools
During manuscript preparation, we used AI assistance for prose shortening and copy-editing, and as coding support for scaffolding evaluation scripts and refactoring utilities. All analysis code, evaluation protocols, and results were designed, implemented, and verified by the authors.
@@ Appendix A A formal presentation of the Erotetic Theory
For $R\in\mathcal{R}$ write $R=\langle U_{R},E_{R},D_{R}\rangle$. For $\langle\emptyset,\emptyset,\emptyset\rangle$ write $0_{\mathcal{R}}$.
@@
$\in\mathbf{T}_{1}$ and $\langle f^{k},\langle t_{1},\ldots,t_{k}\rangle\rangle\in\mathbf{T}_{1}$ whenever $f^{k}\in\mathbf{F}$ and for some $1\leq i\leq k$, $t_{i}\in\mathbf{T}_{1}$ and $t_{j}\in\mathbf{T}$ for $j\neq i$. $\langle P^{k},\langle t_{1},\ldots,t_{k}\rangle\rangle\in\mathcal{A}_{1}$, whenever $P^{k}\in\mathbf{P}$ and for some $1\leq i\leq k$, $t_{j}\in\mathbf{T}$ for all $j\neq i$ and $t_{i}\in\mathbf{T}_{1}$.
@@
$R]_{X}=\langle U_{R}\cap X,E_{R}\cap X,D_{R}\cap((E_{R}\cap X)\times(U_{R}\cap X))\rangle$.
$R]_{\Gamma}=[R]_{\mathbf{A}(\Gamma)}$.
@@
$I]=\{\langle t,x\rangle:\langle t,x\rangle\in I\wedge x[t/?]\in\mathcal{A}(\Gamma\cup\Theta)\}$
@@
$\Gamma^{\Theta}_{RI}[\Delta^{\{0\}}_{SJ}]^{A}=(\arg\max\limits_{\gamma\in\Gamma}\ \Delta[\{\{p\}:p\in\gamma\}]^{\mathcal{A}P})^{\Theta}_{[R][I]}$
@@
$\Gamma^{\Theta}_{RI}]^{N}=(\Theta^{\{0\}}\otimes([\Gamma]^{N})^{\{0\}})_{[R]^{N}[I]^{N}}$
[…]
@@
[…]
$R]^{N}=\langle E_{R},U_{R},\{\langle a,b\rangle\in U_{R}\times E_{R}:\langle b,a\rangle\notin D_{R}\}\rangle$
$I]^{N}=I\cup\{\langle t,\bar{x}\rangle:\langle t,x\rangle\in I\}$
@@
$a\in\mathbf{A}$ is novel during an inference step $\langle C,G\rangle[D]^{O}$ if $a\notin\mathbf{A}(G)\cup\mathbf{A}(D)$ and $a$ has not yet appeared in the computation of the conclusion of the step. A function $\nu:X\to\mathbf{A}$, where $X\subseteq\mathbf{A}$ is finite, is novel during an inference step if each $\nu(x)$ is novel and $\nu$ is injective, and then $\nu]_{X}$ denotes the simultaneous substitution $\nu(a_{1})/a_{1},\ldots,\nu(a_{k})/a_{k}]$, where $X=\{a_{1},\ldots,a_{k}\}$. Let $(-)^{\mathrm{nov}(X)}$ stand for $(-)[\nu]_{X}$ where $\nu:X\to\mathbf{A}$ is novel.
@@
For either $\mathbf{A}(R)\cap\mathbf{A}(S)=\emptyset$ or $R]_{\Delta\cup\Psi}=S$,
$)\Gamma^{\Theta}_{RI}[\Delta^{\Psi}_{SJ}]^{M}=\bigoplus^{R\bowtie S}_{\gamma\in\Gamma}\Big(\{\gamma\}^{\Theta}_{RI}\otimes^{R\bowtie S}\Delta^{\Psi}_{SJ}\otimes^{R\bowtie S}\bigotimes^{R\bowtie S}_{\langle t,u\rangle\in M^{\prime}_{IJ}(\gamma)}\mathrm{Sub}^{R\bowtie S}_{\langle t,u\rangle}(\Delta^{\{0\}}_{J}$$\Big)$,
@@
For $\mathbf{A}(\Gamma)\cap\mathbf{A}(\Theta)=\emptyset$ and either $\mathbf{A}(R)\cap\mathbf{A}(S)=\emptyset$ or $R]_{\Delta}=S$,
[…]
@@
    $\mathbf{A}(\Gamma)\cap\mathbf{A}(\Theta)=\emptyset$ and either $\mathbf{A}(R)\cap\mathbf{A}(S)=\emptyset$ or $R]_{\Delta}=S$.
@@
[…]
@@
(I) If $\mathbf{A}(\Delta)\subseteq\mathbf{A}(\Gamma\cup\Theta)$ and $R]_{\Delta}=S$,